OutbackOutback Escapes generates average revenue of $5,600 per person on its​ 5-day package tours to wildlife parks in Kenya. The variable costs per person are as​ follows:

Annual fixed costs total $560,000

1.

Calculate the number of package tours that must be sold to break even.

2.

Calculate the revenue needed to earn a target operating income of $105,000

3.

If fixed costs increase by $28,000​,what decrease in variable cost per person must be achieved to maintain the breakeven point calculated in requirement​ 1?

4.

The general manager at OutbackOutback Escapes proposes to increase the price of the package tour to $6,200 to decrease the breakeven point in units. Using information in the original​problem, calculate the new breakeven point in units. What factors should the general manager consider before deciding to increase the price of the package​ tour?

Airfare

$1,700

Hotel accommodations

1,500

Meals

500

Ground transportation

300

Park tickets and other costs

200

Total

$4,200
Answer #1

Contribution Margin per person = Revenue per person – Variable costs per person

= 5600-4200

= $1400 per person

CM Ratio = CM/Sales

= 1400/5600

= 25%

1.Break even packages = Fixed costs/Contribution Margin per package

= 560,000/1400

= 400 packages

2.Required Revenue = (105000+560,000)/25%

= 2,660,000

3.Decrease in variable cost per person = 28000/400 = $70 per person

4.Break even point = 560,000/(6200-4200)

= 280 packages

Price charged by competitors

Price elasticity of demand etc.